CALCULATING H_{2}O STATES UP TO DISSOCIATION STATES USING PDVR3D

\underline{HAMSE Y. MUSSA} and JONATHAN TENNYSON, Department of Physics and Astronomy, University College London, Gower St., London WC1E 6BT, UK.

Even small chemically bound molecules have 10^{5} or more bound states. Calculations of this size represent a grand challenge to conventional computers. We have parallelized DVR based program suite DVR3D(1) to give PDVR3D(2). The PDVR3D suite runs on the Cray T3D/T3E at Edinburgh University(U.K.), the IBM SP2 at Daresbury(U.K.) and on the Cray T3E at CINECA (Italy). As a first application of PDVR3D, we are studying the water molecule using two newly available global potentials due to Varandas(3) and Ho and Rabitz(4). We have calculated the ro-vibrational levels of water up to dissociation limits for both potentials. Studies of bound and quasibound ro-vibrational states of H_{3}^{+}, using a realistic global potential(5), are also being performed.

\underline{References:}

[1] J.Tennyson,J.R.Henderson,N.F.Fulton,Computer.Phys.Comm.,86 (1995) 175-196.

[2] H.Y. Mussa, J. Tennyson,C.J. Noble and R.J.Allan, Computer.Phys.Comm.,108 (1998) 29-37.

[3] A.J.C. Varandas, J.Chem.Phys., 105 (1996) 9. [4]T.-S. Ho, T. Hollebeck, H. Rabitz., L.B. Harding, G. Schatz, J.Chem.Phys., 105 (1996) 10472.

[5] R.Prosmiti, O.L. Polyansky and J. Tennyson, Chem. Phys. Lett., 273 (1997) 107-114.